Show that x3+ x2+x+1 is reducible over Q. Does this fact contradict the corollary to Theorem 17.4?


1
Expert's answer
2021-11-26T09:20:34-0500

"x^3+ x^2+x+1=(x^2+1)(x+1)"

since it splits into factors, then it is reducible over Q


corollary:

for any prime p:

"P(x)=\frac{x^p-1}{x-1}=x^{p-1}+x^{p-2}+...+x+1"

is irreducible over Q.


in our case: "x^{p-1}=x^3\implies p=4" is not prime

so, our result does not contradict the corollary